# Judge Audio Captions

### v1

@@ IV Conclusion
This paper introduces $\text{CLAIR}_{A}$, a simple and interpretable domain-specific LLM-based measure for audio captioning. We demonstrate that not only is our simple approach well-aligned with human judgments, but also that such a method is significantly more interpretable to downstream human users. While $\text{CLAIR}_{A}$ is a first step towards LLM evaluation of audio captions, we hope that our work inspired further research into how LLMs can align with human judgment and can be used to develop simple and interpretable systems across a wide range of audio domains.
Acknowledgements: We would like to acknowledge Jeeweon Jung for his helpful contribution to the data used in the paper. GPT-4o was used to check the language of the paper (all sections) for spelling and grammar concerns.

### v2

@@ IV Conclusion
This paper introduces $\text{CLAIR}_{A}$, a simple and interpretable domain-specific LLM-based measure for audio captioning. We demonstrate that not only is our simple approach well-aligned with human judgments, but also that such a method is significantly more interpretable to downstream human users. While $\text{CLAIR}_{A}$ is a first step towards LLM evaluation of audio captions, we hope that our work inspires further research into how LLMs can align with human judgment and can be used to develop simple and interpretable systems across a wide range of audio domains.
Acknowledgements
As part of their affiliation with UC Berkeley, the authors were supported in part by the National Science Foundation, the Ford Foundation, and/or the Berkeley Artificial Intelligence Research (BAIR) Industrial Alliance program. Sky Computing Lab is supported by gifts from Accenture, AMD, Anyscale, Cisco, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Lightspeed, Mibura, Microsoft, NVIDIA, Samsung SDS, and SAP. We thank Jeeweon Jung for his helpful contribution to the data used in the paper. GPT-4o was used to check the language of the paper (all sections) for spelling and grammar concerns.